OpenStudy (anonymous):

38. Tell whether the lines for the pair of equations are parallel, perpendicular, or neither. y = –2/3x + 12x – 3y = –3 (1 point) parallel perpendicular neither

OpenStudy (anonymous):

I don't have many more.

OpenStudy (anonymous):

If you rewrite that second equation you get: y = (1/3)x + 1 Now the slope of the first equation is -2/3 and the slope of the second equation is 1/3 Since the slopes are not equal, nor are they negative reciprocals of each other, they are not parallel or perpendicular, so the answer is : neither

OpenStudy (anonymous):

Sound good, @sgstiles ?

OpenStudy (anonymous):

I have 4 more I think.

OpenStudy (anonymous):

37. Write y = –2/5x – 3 in standard form using integers. (1 point)2x + 5y = –152x + y = –32x + 5y = –3x + y = –15

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @sgstiles

OpenStudy (anonymous):

ok, just close this post and open a new one and I'll see you there!